Factor each polynomial by splitting the middle term and grouping

6x squared plus x minus 12 = the answer should be (2x plus 3)(3X MINUS 4)
but I can't seem to get it there/ please help/thanks

Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

multiply the factored form.
(2x+3)(3x-4)= 6x^2+x-12
take (2x)(3x)=6x^2
(2x)(-4)+(3)(3x)=x
(3)(-4)=-12

Answer by checkley77(12844)

6x^2+x-12
The factors of 6 are (1 & 6), (-1 & -6), (2 & 3), (-2 &-3).
The factors of 12 are (-1 & 12), (1 & -12), (-2 & 6), (2 & -6), (-3 & 4),
(3 & -4).
Now you need to select a set from each term that when multiplied together will equal +1, the middle term.
Examples:
Use the (1,6) & (-1,12)
(1*-1)+(6*12)=-1*72=71 not close.
Use the (2,3) & (-2,6)
(2*-2)+(3*6)=-4+18=14 closer.
Use the (2,3) & (-4,3)
(2*-4)+(3*3)=-8+9=1 looks like a winner.
Thus we have: (2x+3)(3x-4)using FOIL we get.
2x*3x=6x^2
2x*-4=-8x
3*3x=9x
3*-4=-12
6x^2-8x+9x-12
6x^2+x-12